1. Find the measure of the larger of two supplementary angles if the larger is 4 times the smaller.
Answer by rapaljer(4671)

Let x= smaller
4x= larger

x+4x=180
5x=180
x=180/5=36 degrees = smaller angle
4x= 144 degrees = larger angle

Check: 36 + 144 = 180

Dr. Robert J. Rapalje, Retired
Seminole State College of Florida
Altamonte Springs Campus